# A linear-time algorithm computing the resident fitness in interacting trajectories

### v1

@@ 3 Results
For a sequence $(\beth^{(n)})_{n\in\mathbb{N}}$ of non-degenerate initial data $\beth^{(n)}=((t_{i}^{(n)},a_{i}^{(n)}))_{1\leq i<n+1}$, writing $\mathbb{H}(\beth^{(n)})=(h_{i}^{(n)})_{0\leq i<n+1}$, we denote by
$k_{n}(\beth^{(n)})$ $=\sum_{t\geq 0\colon t\text{ resident change time of }\mathbb{H}(\beth^{(n)})}\sum_{i=0}^{n}\mathds{1}_{\{h_{i}^{\prime(n)}(t-)\neq h_{i}^{\prime(n)}(t+)\}\cap\{h_{i}^{(n)}(t)>0\}}$
$=\sum_{t\geq 0\colon t\text{ resident change time of }\mathbb{H}(\beth^{(n)})}\big{|}\{i\in\{0,1,\ldots,n\}\colon h_{i}\text{ is alive at time }t\}\big{|}$
@@
The proof of Proposition 9 can be found in Section 4.1.
Remark 10
@@
Apart from the slope changes corresponding to the definition of $k_{n}$, at each time $t_{i}$, $i=1,\ldots,n$, there is an additional slope change where the slope of the $i$-th trajectory switches from 00 to $a_{i}$. Moreover, at most $n$ trajectories eventually hit 0 again and then their slope drops to 0. Altogether, these are at most $2n$ additional slope changes, and precisely $2n$ unless there are multiple indices $j$ with the same value of $f(t_{j})+a_{j}$ as the ultimate resident. (In the case of the PIT, the latter is only possible if $\mathbb{P}(A_{i}=c)>0$ holds for some $c>0$.) Hence, given Proposition 9, ignoring these $O(n)$ additional slope changes will not yield substantial changes in the worst-case estimates of the number of kinks.
Given Proposition 9, computing the slopes of all trajectories in a system of $n$ interacting trajectories clearly takes $\Omega(n^{2})$ time. One might think that even computing the resident fitness for all times takes a quadratic amount of time. However, we have the following theorem, where $\beth$ counts as ordered if the corresponding $t_{i}$’s are sorted in the input and unordered otherwise.
Theorem 11
@@
The algorithmic complexity of finding $t\mapsto(\varrho^{(\beth)}(t))_{t\geq 0}$ for any finite, nondegenerate, ordered $\beth$ of size $n$ is $\Theta(n)$, and it is $\Theta(n\log n)$ in the unordered case.
@@
An algorithm proving the theorem will be provided in Section 4.2. The key idea is to use the continued lines representation, which makes it possible to provide such an algorithm using a simple data structure (involving two pointers and a doubly linked list), despite the possibly large number of kinks.
The output of the algorithm is a list containing the time intervals of residencies in consecutive order, together with the indices of the residents and the corresponding resident fitness values. Given this output, in Section 5.1 we explain how to determine the set of trajectories that reach fixation and the ancestral relations between these trajectories (the notion of fixation will also be explained there), whereas in Section 5.2 below we summarize how to compute the resident fitness at a given time and the trajectory $(h_{i}(t))_{t\geq 0}$ for fixed $i$ efficiently.
@@
Our next proposition shows that if one considers the PIT where the birth times of trajectories form a Poisson process, then even the expected number of kinks will only grow linearly in time. (Due to the time-continuous nature of the Poisson process, here one has to work with a notion of kinks (3) slightly different from the one (3) above; see Remark 13 for a related discussion.)
@@
Let $\lambda,\gamma$ as above and let us consider the $\mathrm{PIT}(\lambda,\gamma)$. For $t\geq 0$ we write $N_{t}=\{i\geq 1\colon T_{i}\leq t\}$ where $T_{i}$ is the birth time of the $i$-th trajectory (with the aforementioned convention that $T_{0}=0$), then $N_{t}$ is $\mathrm{Poisson}(\lambda t)$ distributed for any $t>0$, and for any $n\in\mathbb{N}$ and $0=t_{0}<t_{1}<\ldots<t_{n}$, $N_{t_{i}}-N_{t_{i-1}}$, $i=1,\ldots,n$ are independent). Define the total number of kinks up to $t$ as
@@
$K_{t}=\sum_{0<s\leq t\colon s\text{ resident change time}}\ \sum_{i\in\mathbb{N}_{0}}\mathds{1}_{\{H_{i}^{\prime}(t-)\neq H_{i}^{\prime}(t+)\}}.$
Proposition 12
@@
The proof of Proposition 12 will be carried out in Section 4.3. This proof is based on an extension of the aforementioned renewal argument of [13], combined with classical large-deviation inequalities. The main message of this proof is that the clustering of trajectories that leads to many kinks (like in the construction in the proof of Proposition 9) is exponentially unlikely in the number of participating trajectories if these trajectories follow each other directly in the PIT according to their birth times.
Remark 13 (Relation between Propositions 9 and 12)
@@
the difference $k_{N_{t}}(\beth^{(N_{t})})-K_{t}$ being the number of those kinks in $\mathbb{H}(\beth^{(N_{t})})$ that happen after time $t$. However, it will be apparent from the proof of Proposition 12 that following time 0, after a stopping time with finite expectation, a state will be reached where $H_{i}(t)=0$ for all but one $i$. Hence, by the strong Markov property, this state will be visited infinitely often, with i.i.d. waiting times with finite expectation inbetween. Since a trajectory can only have kinks before its height ultimately reaches 00, it follows that the difference between $K_{t}$ and $k_{N_{t}}(\beth^{(N_{t})})$ is of finite order. This way, Proposition 12 can indeed be seen as complementary to Proposition 9 in the Poissonian case.
@@ 4 Proofs
4.1 Lower bound construction for the number of kinks: proof of Proposition 9
Proof: For $n\in\mathbb{N}$ describe an input $\beth^{(n)}$ so that $(\mathbb{H}(\beth^{(n)})_{n\in\mathbb{N}}$ satisfies the conditions of the proposition and we verify that it indeed satisfies them. For simplicity, we use the continued lines representation introduced in Remark 8 (which is in bijection with the representation as a system of interacting trajectories).
@@ 4.1 Lower bound construction for the number of kinks: proof of Proposition 9
$l_{0}\text{ is resident on }[0,1)\text{ and for $k\geq 1$, }l_{k}\text{ is resident on }\big{[}1+\frac{2(k-1)}{n^{2}},1+\frac{2k}{n^{2}}\big{]}.$ (10)
Indeed, recall that at a given time $t$, the index of the resident is the index of the ray that is at the highest position among all rays at time $t$ (which implies that the ray has already been started by time $t$, and in the case of a draw, the index with the higher slope is the index of the resident). Now, (10) follows inductively from the fact that 00 is obviously resident at time 00, type $k$ cannot be resident at time $t<1+\frac{2(k-1)}{n^{2}}$ because then $l_{k-1}(t)>l_{k}(t)$, and it cannot be resident at time $t>1+\frac{2k}{n^{2}}$ either because then $l_{k+1}(t)>l_{k}(t)$, but at any time there must be a resident type, so $l_{k}$ must be resident for $\frac{2(k-1)}{n^{2}}<t<\frac{2k}{n^{2}}$ and thus by construction also for $t=\frac{2(k-1)}{n^{2}}$.
Now we claim that the 00-th trajectory of $\mathbb{H}(\beth^{(n)})$ is still alive at time $1+\frac{2(n-1)}{n^{2}}$, equivalently, $l_{0}(1+\frac{2(n-1)}{n^{2}})>l^{r}(1+\frac{2(n-1)}{n^{2}})-1=l_{n}(1+\frac{2(n-1)}{n^{2}})-1$. Indeed, we have that
@@
By construction, for $t>1+\frac{2(n-1)}{n^{2}}$, $l_{0}(t)$ is definitely the smallest one among all values $l_{i}(t)$, $i\in\{0,1,\ldots,n\}$. This implies that all other trajectories of $\mathbb{H}(\beth^{(n)})$ are also alive at time $1+\frac{2(n-1)}{n^{2}}$. Since they are all already born (i.e. the corresponding rays of the continued lines representation are already started) before time $1$, it follows that at all resident change times all trajectories of $\mathbb{H}(\beth^{(n)})$ are alive and thus they change their slopes. This yields $k_{n}(\beth^{(n)})=n(n+1)=\Omega(n^{2})$.  $\Box$
4.2 Proof of Theorem 11
The algorithm for the ordered case processes births one at a time, in ascending order, while maintaining a list of residents. For trajectory $i$, the parent is found by linear search from the start of the resident list forwards, and whether it becomes resident is determined by linear search from the end of the resident list backwards.
Recall from Remark 8 that in the continued lines representation, each trajectory birth is a pair $(t_{i},a_{i})$, where $t_{i}$ is the birth time and $a_{i}$ is the slope relative to its parent.
A trajectory is a tuple $(t_{i},p_{i},b_{i},y_{i})$, where $t_{i}$ is the birth time, $p_{i}$ is the index of the trajectory of the parent, who is the resident at time $t_{i}$, $b_{i}=b_{p_{i}}+a_{i}$ is the slope, and $y_{i}$ is the initial height, which is one less than the height of the parent at time $t_{i}$. This corresponds to a ray starting at point $(t_{i},y_{i})$ with slope $b_{i}$. A residency is a pair $(s_{j},m_{j})$, where $s_{j}$ is the residency start time and $m_{j}$ is the trajectory index. A residency starts when its trajectory intersects the trajectory of the current resident, so the residencies form a convex broken line. The $j$th residency starts at point $(s_{j},y_{m_{j}}+b_{m_{j}}(s_{j}-t_{m_{j}}))$.
Initially, there is one trajectory in the system, $(t_{0}=0,p_{0}=-1,b_{0}=0,y_{0}=1)$, which is the current resident, $(s_{0}=0,m_{0}=0)$. In step $i$, the forward linear search finds the parent of trajectory $i$, which is the last resident $j$ with $s_{j}\leq t_{i}$. The $i$th trajectory is $(t_{i},p_{i}=m_{j},b_{i}=(b_{p_{i}}+a_{i}),y_{i}=(y_{p_{i}}+b_{p_{i}}(t_{i}-t_{p_{i}})-1))$. Since the input is ordered by $t_{i}$, the search can resume at the position it stopped at in the previous step.
@@ 4.2 Proof of Theorem 11
The list of residents may be stored in a doubly linked list. Since both linear searches pass each trajectory at most once, the algorithm has time $O(n)$. It is easy to see, that it is also $\Theta(n)$, for example when all trajectories become residents.
When the input is not ordered, it can be shown that the optimal time for solving the problem is $\Omega(n\log n)$, since for a given unordered list of pairwise distinct integers $\{x_{1},\dots,x_{n}\}$, the input trajectories $\{(x_{1},2),\dots,(x_{n},2)\}$ will result in a resident list containing all $x_{i}$, in sorted order. It is also $\Theta(n\log n)$ because the same algorithm with initial sorting can be used.
4.3 Proof of Proposition 12
Proof: Let $E_{0}=0<E_{1}<E_{2}<\ldots$ be the renewal extinction times, defined via
$E_{i}=\inf\{t\geq E_{i-1}\mid\text{ only the resident type at time $t$ has a positive height}\}$
@@ 4.3 Proof of Proposition 12
for $i\geq 1$. Note that for $i\geq 1$, on the time interval $(E_{i-1},E_{i}]$ only trajectories born in $(E_{i-1},E_{i}]$ can become resident, and trajectories whose height function can suffer any kinks on this time interval due to resident changes are also the same trajectories plus the trajectory that was resident at time $E_{i-1}$. Indeed, any other trajectory born on $[0,E_{i-1}]$ goes extinct by the time $E_{i-1}$. Hence, for $t\geq 0$, putting $\eta_{t}=|\{i\geq 1\colon E_{i}\leq t\}|=\sup\{i\geq 0\colon E_{i}\leq t\}$, we have $E_{\eta_{t}}\leq t<E_{\eta_{t}+1}$. Note also that $\eta_{t}\leq N_{t}$ for all $t$ because at each $E_{i}$ a trajectory born before $E_{i}$ becomes resident (and one trajectory can only become resident at most once).
@@
It can be proven analogously to the proof of [13, Lemma 5.5] that $E_{i}-E_{i-1}$ is stochastically dominated by a geometric random variable, and the random variables $(E_{i}-E_{i-1})_{i\geq 1}$ are i.i.d. (in other words, $(E_{n})_{n\in\mathbb{N}}$ is a renewal process). They have a finite first moment $m_{1}:=\mathbb{E}(E_{1})$ and a finite second moment $m_{2}:=\mathbb{E}(E_{1}^{2})$. This together with the tower property implies that
@@
$=\mathbb{E}(\mathbb{E}((N_{E_{i}}-N_{E_{i-1}})(N_{E_{i}}-N_{E_{i-1}}+1)|E_{i}-E_{i-1}))=\mathbb{E}(\lambda^{2}(E_{i}-E_{i-1})^{2}+2\lambda(E_{i}-E_{i-1}))$
$=\lambda^{2}m_{2}+2\lambda m_{1}.$
@@
almost surely (cf. [9, Theorem 2.5.10] for the latter assertion). Using Hölder’s inequality and (4.3), we conclude that for $\delta>0$, for all $t$ sufficiently large

### v2

@@ 3 Results
For a sequence $(\beth^{(n)})_{n\in\mathbb{N}}$ of nondegenerate initial data $\beth^{(n)}=((t_{i}^{(n)},a_{i}^{(n)}))_{1\leq i<n+1}$, writing $\mathbb{H}(\beth^{(n)})=(h_{i}^{(n)})_{0\leq i<n+1}$, we denote by
$k_{n}(\beth^{(n)})$ $=\sum_{t\geq 0\colon t\text{ resident change time of }\mathbb{H}(\beth^{(n)})}\sum_{i=0}^{n}\mathds{1}_{\{v_{i}^{(n)}(t-)\neq v_{i}^{(n)}(t+)\}\cap\{h_{i}^{(n)}(t)>0\}}$
[…]
@@
The proof of Proposition 3.1 can be found in Section 4.1.
Remark 3.2.
@@
Apart from the slope changes corresponding to the definition of $k_{n}$, at each time $t_{i}$, $i=1,\ldots,n$, there is an additional slope change where the slope of the $i$-th trajectory switches from 0 to $a_{i}$. Moreover, at most $n$ trajectories eventually hit 0 again and then their slope drops to 0. Altogether, these are at most $2n$ additional slope changes, and precisely $2n$ unless there are multiple indices $j$ with the same value of $f(t_{j})+a_{j}$ as the ultimate resident. (In the case of the PIT, the latter is only possible if $\mathbb{P}(A_{1}=c)>0$ holds for some $c>0$.) Hence, given Proposition 3.1, ignoring these $O(n)$ additional slope changes will not yield substantial changes in the worst-case estimates of the number of kinks.
[…]
Theorem 3.3.
@@
The algorithmic complexity of finding $t\mapsto\varrho^{(\beth)}(t)$, $t\geq 0$, for any finite, nondegenerate, $red \beth o$f size $n$ is $\Theta(n)$.
@@
The proof of this theorem will be provided in Section 4.2. The proof of the upper bound is constructive: we present an algorithm that computes the array describing the function in $O(n)$ time. The key idea is to use the continued lines representation of the system, which makes it possible to provide such an algorithm using a simple data structure, despite the possibly large number of kinks. As a byproduct, the algorithm also computes the trajectories in the continued lines representation of the system. From these outputs, the resident fitness function $t\mapsto f(t)$ is also easily computable, since it corresponds to the slopes of the trajectories of the residents.
[…]
In the special case when we additionally know that each $a_{i}$ is contained in the set $[k]$ for some $k\geq 1$, the algorithm appearing in the proof of Theorem 3.3 can easily be modified so that it also computes the piecewise slopes of all interacting trajectories (not only the resident fitness or the slopes of the rays of the continued lines representation) in $O(kn)$ time. This fact is reflected by the following corollary.
Corollary 3.4.
Let $k\in\mathbb{N}$. For any finite, nondegenerate, $red \beth o$f size $n$ such that each initial slope $a_{i}$ lies in $[k]$, the family of piecewise constant slopes $(t\mapsto v_{i}(t))_{i\in[n]}$ can be computed via a deterministic algorithm of runtime $O(kn)$.
The proof of Corollary 3.4 can be found in Section 4.3.
@@
Our next proposition shows that if one considers the PIT where the birth times of trajectories form a Poisson process, then even the expected number of kinks will only grow linearly in time. Due to the time-continuous nature of the Poisson process, here one has to work with a notion of kinks (3) slightly different from the one in (3) above; see Remark 3.6 below for a related discussion.
@@
Let $\lambda,\gamma$ as above and let us consider the $\mathrm{PIT}(\lambda,\gamma)$. For $t\geq 0$ we write $N_{t}=|\{i\geq 1\colon T_{i}\leq t\}|$ where $T_{i}$ is the birth time of the $i$-th trajectory (with the aforementioned convention that $T_{0}=0$), then $N_{t}$ is $\mathrm{Poisson}(\lambda t)$ distributed for any $t>0$, and for any $n\in\mathbb{N}$ and $0=t_{0}<t_{1}<\ldots<t_{n}$, $N_{t_{i}}-N_{t_{i-1}}$, $i=1,\ldots,n$ are independent. Define the total number of kinks up to $t$ as
@@
$K_{t}=\sum_{0<s\leq t\colon s\text{ resident change time}}\ \sum_{i\in\mathbb{N}_{0}}\mathds{1}_{\{H_{i}^{\prime}(t-)\neq H_{i}^{\prime}(t+)\}\cap\{H_{i}(t)>0\}}.$
Proposition 3.5.
@@
The proof of Proposition 3.5 will be carried out in Section 4.4. This proof is based on an extension of the aforementioned renewal argument of [16], combined with classical large-deviation inequalities. The main message of this proof is that the clustering of trajectories that leads to many kinks (like in the construction in the proof of Proposition 3.1) is exponentially unlikely in the number of participating trajectories if these trajectories follow each other directly in the PIT according to their birth times.
Remark 3.6 (Relation between Propositions 3.1 and 3.5).
@@
the difference $k_{N_{t}}(\beth^{(N_{t})})-K_{t}$ being the number of those kinks in $\mathbb{H}(\beth^{(N_{t})})$ that happen after time $t$. However, it will be apparent from the proof of Proposition 3.5 that following time 0, after a stopping time with finite expectation, a state will be reached where $H_{i}(t)=0$ for all but one $i$. Hence, by the strong Markov property, this state will be visited infinitely often, with i.i.d. waiting times of finite expectation inbetween. Since a trajectory can only have kinks before its height ultimately reaches 0, it follows that the difference between $K_{t}$ and $k_{N_{t}}(\beth^{(N_{t})})$ is of finite order. This way, Proposition 3.5 can indeed be seen as complementary to Proposition 3.1 in the Poissonian case.
Based on the proof of Proposition 3.5 and the one of [16, Theorem 2.8], in Section 4.4 we will also derive the following corollary, which can be seen as a strong law of large numbers for the total number of kinks.
Corollary 3.7.
Let $E_{0}=0<E_{1}<E_{2}<\ldots$ be the renewal extinction times, defined via
[…]
for $i\geq 1$. We have
$\lim_{t\to\infty}\frac{K_{t}}{t}=\frac{\mathbb{E}(K_{E_{1}})}{\mathbb{E}(E_{1})},$
almost surely. The limit on the right-hand side is in $(0,\infty)$ for any choice of $\gamma$.
Remark 3.8.
Corollary 3.7 is formally analogous to [16, Theorem 2.8] on the speed of adaptation: There, under the additional assumption that $\mathbb{E}(A_{1})<\infty$, using the solitary resident change times $L_{0}=0<L_{1}<L_{2}<\ldots$ defined as
[…]
it was shown that the resident fitness $t\mapsto F(t)$ satisfies
$\lim_{t\to\infty}\frac{F(t)}{t}=\frac{\mathbb{E}(F(L_{1}))}{\mathbb{E}(L_{1})}\in(0,\infty),$
almost surely. We will explain this analogy in more detail and exploit it in the proof of Corollary 3.7 below. Note however that a substantial difference between the two results is that if $\mathbb{E}(A_{1})=\infty$, then $\lim_{t\to\infty}F(t)/t=\infty$ a.s., while in Corollary 3.7 the limit is always finite.
Remark 3.9.
When $A_{1}$ is constant and equal to $c>0$, the limit in (3.7) is $2\lambda$. To see this, note that if all $A_{i}$’s are equal to $c$, then any trajectory born after time 0 will start with slope $c$ at height 0, reach slope 0 at the first resident change following its birth and reach slope c-c at the next resident change. This slope will be preserved until the trajectory reaches height 0 (see also Lemma 4.1). Now, by time $E_{1}$, all trajectories born before $E_{1}$ but the one that is resident at time $E_{1}$ must have reached height 0 again. This accounts for 2 slope changes due to resident changes per trajectory, apart from the 0-th trajectory, which started from height 1 with slope 0, and the trajectory that is resident at time $E_{1}$, which ends at height 1 with slope 0; the latter two trajectories have only had one slope change by time $E_{1}$ each. Thus, almost surely, $\mathbb{E}(K_{E_{1}}|N_{E_{1}})=2N_{E_{1}}$, and hence by the tower property, $\mathbb{E}(K_{E_{1}})=2\mathbb{E}(N_{E_{1}}|N_{E_{1}})=2\mathbb{E}(N_{E_{1}})=2\mathbb{E}(\mathbb{E}(N_{E_{1}}|E_{1}))=2\lambda\mathbb{E}(E_{1})$, which implies that the limit in  (3.7) is indeed $2\lambda$. For general distributions $\gamma$, we expect that computing the limit in (3.8) is harder.
As we already anticipated in Remark 2.7[16] studied a variant of the PIT where trajectories have a chance to stay constant equal to zero, which corresponds to rapid extinction of the given mutant subpopulation in the Moran model. There, in the constant-$c$ case, trajectories arrive at rate $\lambda$ but independently with probability $\frac{1}{1+c}$ they stay constant 0 and only with probability $\widetilde{\lambda}:=\frac{\lambda c}{1+c}$ they follow the dynamics described in Section 1. In that setting, $\lim_{t\to\infty}K_{t}/t$ would be $2\widetilde{\lambda}=\frac{2\lambda c}{1+c}$. We note that this is also the only case where an explicit formula for the speed of adaptation is known. We have (in the latter variant of the PIT) that $\lim_{t\to\infty}F(t)/t=\frac{\lambda c^{2}}{1+c+\lambda}$. This result is [16, Theorem 2.10], which in turn originates from [1, Section 3.1]; see Section 5.2 for more details.
@@ 4 Proofs
4.1 Lower bound construction for the number of kinks: proof of Proposition 3.1
Proof.
The sequence $(\mathbb{H}(\beth^{(n)})_{n\in\mathbb{N}}$ corresponding to the statement of the proposition is given as follows. For simplicity, we use the continued lines representation introduced in Remark 2.8 (which is in bijection with the representation as a system of interacting trajectories).
@@
[…]
Indeed, recall that at a given time $t$, the index of the resident is the index of the ray that is at the highest position among all rays at time $t$ (which implies that the ray has already been started by time $t$, and in the case of a draw, the index with the higher slope is the index of the resident). Now, (13) follows inductively from the fact that 0 is obviously resident at time 0, type $k$ cannot be resident at time $t<1+\frac{2(k-1)}{n^{2}}$ because then $l_{k-1}(t)>l_{k}(t)$, and it cannot be resident at time $t>1+\frac{2k}{n^{2}}$ either because then $l_{k+1}(t)>l_{k}(t)$, but at any time there must be a resident type, so $l_{k}$ must be resident for $\frac{2(k-1)}{n^{2}}<t<\frac{2k}{n^{2}}$ and thus by construction also for $t=\frac{2(k-1)}{n^{2}}$.
Now we claim that the 0-th trajectory of $\mathbb{H}(\beth^{(n)})$ is still alive at time $1+\frac{2(n-1)}{n^{2}}$, equivalently, $l_{0}(1+\frac{2(n-1)}{n^{2}})>l^{r}(1+\frac{2(n-1)}{n^{2}})-1=l_{n}(1+\frac{2(n-1)}{n^{2}})-1$, where we recall from Remark 2.8 that $l^{r}(t)$ is the height of the current resident trajectory at time $t$. Indeed, we have that
@@
By construction, for $t>1+\frac{2(n-1)}{n^{2}}$, $l_{0}(t)$ is definitely the smallest one among all values $l_{i}(t)$, $i\in\{0,1,\ldots,n\}$. This implies that all other trajectories of $\mathbb{H}(\beth^{(n)})$ are also alive at time $1+\frac{2(n-1)}{n^{2}}$. Since they are all already born (i.e. the corresponding rays of the continued lines representation are already started) before time $1$, it follows that at all resident change times all trajectories of $\mathbb{H}(\beth^{(n)})$ are alive and thus they change their slopes. This yields $k_{n}(\beth^{(n)})=n(n+1)=\Omega(n^{2})$. ∎
4.2 Proof of Theorem 3.3
The algorithm processes births one at a time, in increasing order, while maintaining an ordered list of residents. For trajectory $i$, the parent is found by linear search from the start of the resident list forwards, and whether it becomes resident is determined by linear search from the end of the resident list backwards.
Recall from Remark 2.8 that in the continued lines representation, each trajectory birth is a pair $(t_{i},a_{i})$, where $t_{i}$ is the birth time and $a_{i}$ is the slope relative to its parent.
A trajectory is a tuple $(t_{i},p_{i},b_{i},y_{i})$, where $t_{i}$ is the birth time, $p_{i}$ is the index of the trajectory of the parent, who is the resident at time $t_{i}$, $b_{i}=b_{p_{i}}+a_{i}$ is the slope, and $y_{i}$ is the initial height, which is one less than the height of the parent at time $t_{i}$. This corresponds to a ray starting at point $(t_{i},y_{i})$ with slope $b_{i}$. A residency is a pair $(s_{j},m_{j})$, where $s_{j}$ is the residency start time and $m_{j}$ is the trajectory index. A residency starts when its trajectory intersects the trajectory of the current resident from below, so the residencies form a convex broken line. The $j$-th residency starts at point $(s_{j},y_{m_{j}}+b_{m_{j}}(s_{j}-t_{m_{j}}))$.
Initially, there is one trajectory in the system, $(t_{0}=0,p_{0}=-1,b_{0}=0,y_{0}=1)$, which is the current resident, $(s_{0}=0,m_{0}=0)$. When processing the $i$-th birth, the forward linear search finds the parent of trajectory $i$ in the list of residents, which is the last resident $j$ with $s_{j}\leq t_{i}$. The $i$-th trajectory is $(t_{i},p_{i}=m_{j},b_{i}=(b_{p_{i}}+a_{i}),y_{i}=(y_{p_{i}}+b_{p_{i}}(t_{i}-t_{p_{i}})-1))$. Since the input is ordered by $t_{i}$, the search for the parent can start from the parent of the previous trajectory.
@@
Since both linear searches pass each trajectory at most once, the algorithm has runtime $O(n)$. Note here that as a byproduct, this algorithm also computed the trajectories $(t_{i},p_{i},b_{i},y_{i})$ in the continued lines representation of the system.
The lower bound construction appearing in the proof of Proposition 3.1 (see Section 4.1) shows that for all $n\in\mathbb{N}$ there exists $\beth^{(n)}$ such that $\mathbb{H}(\beth^{(n)})$ exhibits $n$ resident changes. Therefore, the size of the output of any algorithm computing $t\mapsto\varrho^{(n)}(t)$ is $\Omega(n)$ in the worst case, and hence the runtime of any such algorithm is also $\Omega(n)$.
We conclude that the runtime complexity is indeed $\Theta(n)$, which completes the proof of Theorem 3.3.
4.3 Proof of Corollary 3.4
The key step of the proof of Corollary 3.4 is the following lemma.
Lemma 4.1.
[…]
Proof.
Throughout the proof, we will use the following slightly generalized version of [16, Lemma 2.5] (whose proof is analogous to the one of the original lemma). For all $i\neq j$, recalling that
$e_{i}=\inf\{t>t_{i}\colon h_{i}(t)=0\},$
i.e. $e_{i}$ is the first time when the trajectory $h_{i}$ hits 0 again after its birth time $t_{i}$, we have that
$t\mapsto v_{i}(t)-v_{j}(t)\text{ is constant on }[\max\{t_{i},t_{j}\},\min\{e_{i},e_{j}\}).$
Now assume for a contradiction that for some $t\geq 0$ and $i\geq 0$ we have $v_{i}(t)\leq-2k$. Put
$\tau=\inf\{s\geq 0\colon v_{i}(s)\leq-2k\}.$
It is clear that our assumption is equivalent to the assumption that $\tau<\infty$. Recall that $h_{i}$ is continuous and piecewise linear and hence its right derivative $v_{i}$ is right-continuous. Thus, if $\tau<\infty$, then $v_{i}(\tau)\leq-2k$ holds, and thanks to the non-degeneracy assumption, there is a unique $j\in\mathbb{N}_{0}$ such that the $j$-th trajectory becomes resident at time $\tau$.
Using that $v_{j}(s)\leq k$ for all $s$ and that the $j$-th trajectory becomes resident at time $\tau$, it follows that
$\tau\geq t_{j}+\frac{1}{k}.$
Moreover, by construction, under the assumption $\tau<\infty$, $v_{i}(t_{j})-v_{i}(\tau)$ equals the increment of the resident fitness between times $t_{j}$ and $\tau$ (see (4.3) or also [16, Lemma 2.5]), which is precisely equal to $a_{j}\leq k$. It follows that $v_{i}(t_{j})\leq-k$. Thus, if $\tau<\infty$, then $v_{i}(s)\leq-k$ for all $s\in[t_{j},\tau)$. Therefore, using (4.3) we have
$h_{i}(\tau)=h_{i}(t_{j})+\int_{t_{j}}^{\tau}v_{i}(s)\mathrm{d}s<1-k(\tau-t_{j})<1-k\cdot\frac{1}{k}<0.$
This is a contradiction, and hence it follows that $h_{i}$ hits 0 before time $\tau$, i.e. $e_{i}<\tau$ and thus $\tau=\infty$. This finishes the proof of the lemma. ∎
Now we can carry out the proof of Corollary 3.4.
Proof of Corollary 3.4.
Note that if each $a_{i}$ is integer-valued, then so is $v_{i}(t)$ for all $0\leq i<n+1$ and $t\geq 0$. Since at any resident change time, the slope of any trajectory that is currently at a positive height decreases by at least one and, thanks to Lemma 4.1, the slope of any trajectory at any time ranges between (2k1)-(2k-1) and $k$, it follows that the slope of any trajectory can change at most 3k13k-1 times while the trajectory is at a strictly positive height. In Section 5.3 we describe how to compute the height functions $t\mapsto h_{i}(t)$ for all trajectories in $O(kn)$ time, assuming each trajectory undergoes at most $O(k)$ kinks, which completes the proof of this corollary. ∎
Remark 4.2.
[…]
$t_{2}+\frac{1}{k}<t_{1}+\frac{2}{k-1}-\frac{1}{k}+\frac{1}{k}=t_{1}+\frac{2}{k-1},$
until time $t_{2}+\frac{1}{k}$ the 0-th trajectory does not yet reach height zero again. Consequently, at this time, its slope drops by $v_{2}(t_{2}+\frac{1}{k}-)=k$, i.e. it becomes (2k1)-(2k-1).
In the case of the PIT, if $\lambda>0$ and the support of the measure $\gamma$ contains both $k$ and k1k-1 (i.e. $\mathbb{P}(A_{i}=k)$ and $\mathbb{P}(A_{i}=k-1)$ are both positive), the event that the scenario described in the previous paragraph occurs with $(t_{i},a_{i})=(T_{i},A_{i})$ and $h_{i}=H_{i}$, $i=1,2,3$, has positive probability. Using a restart argument based on e.g. the renewal extinction times defined below in Section 4.4, it follows that for such $\gamma$, the $\mathrm{PIT}(\lambda,\gamma)$ exhibits infinitely many trajectories whose slope reaches (2k1)-(2k-1), almost surely. The same is clearly false e.g. if $k$ is even and the support of $\gamma$ is contained in $\{2,4,\ldots,k\}$.
While Lemma 4.1 guarantees that trajectories can never have slopes less than or equal to 2k-2k when all initial slopes $a_{i}$ are in $(0,k]$, slopes in ((2k1),2k)(-(2k-1),-2k) can be obtained for certain choices of the possible values of $a_{i}$ (using an analogue of the construction in the first paragraph of the present remark). Moreover, since the slope decrement due to one kink can be arbitrarily small, having a finite lower bound on the minimal slope does not imply in general that the number of kinks of a given trajectory is bounded by a constant times the absolute value of the minimum possible slope. This is an obstacle to generalizing Corollary 3.4 to this case.
An analogue of Corollary 3.4 holds whenever the possible initial slopes $a_{i}$ are in $\{\alpha,2\alpha,\ldots,k\alpha\}$ for some $\alpha>0$.
4.4 Proof of Proposition 3.5 and Corollary 3.7
Proof of Proposition 3.5.
@@
Recall the renewal extinction times $E_{i}$ from (3.7). Note that for $i\geq 1$, on the time interval $(E_{i-1},E_{i}]$ only trajectories born in $(E_{i-1},E_{i}]$ can become resident, and only the height functions of the same trajectories plus the one of the trajectory that was resident at time $E_{i-1}$ can suffer any kinks on this time interval due to resident changes. Indeed, any other trajectory born on $0,E_{i-1}]$ goes extinct by the time $E_{i-1}$. Hence, we have $K_{E_{i}}-K_{E_{i-1}}\leq(N_{E_{i}}-N_{E_{i-1}})(N_{E_{i}}-N_{E_{i-1}}+1)$, and for $t\geq 0$, putting $\eta_{t}=|\{i\geq 1\colon E_{i}\leq t\}|=\sup\{i\geq 0\colon E_{i}\leq t\}$, we have $E_{\eta_{t}}\leq t<E_{\eta_{t}+1}$. Note also that $\eta_{t}\leq N_{t}$ for all $t$ because at each $E_{i}$ a trajectory born before $E_{i}$ becomes resident (and one trajectory can only become resident at most once).
@@
It can be proven analogously to the proof of [16, Lemma 5.2] that $E_{i}-E_{i-1}$ is stochastically dominated by a geometric random variable, and the random variables $(E_{i}-E_{i-1})_{i\geq 1}$ are i.i.d. (in other words, $(E_{n})_{n\in\mathbb{N}}$ is a renewal process). They have a finite first moment $m_{1}:=\mathbb{E}(E_{1})$ and a finite second moment $m_{2}:=\mathbb{E}(E_{1}^{2})$. This together with the tower property implies that
@@
$=\mathbb{E}(\mathbb{E}((N_{E_{i}}-N_{E_{i-1}})(N_{E_{i}}-N_{E_{i-1}}+1)|E_{i}-E_{i-1}))$
$=\mathbb{E}(\lambda^{2}(E_{i}-E_{i-1})^{2}+2\lambda(E_{i}-E_{i-1}))=\lambda^{2}m_{2}+2\lambda m_{1}.$
@@
almost surely (cf. [11, Theorem 2.5.10] for the latter assertion). Using Hölder’s inequality, (4.4) and and (4.4), we conclude that for $\delta>0$, for all $t$ sufficiently large